QUESTION 1 50 MARKS
You have recently joined Slam (Pty) Ltd, a company that manufactures and distributes brake pads to the automotive industry, as a financial accountant. The managing director and majority shareholder has asked you to assist him in interpreting the draft financial results for the year ended 30 June 2020 and to review the budget for the new financial year. The company uses first-in-fist out method of inventory valuation. As part of his preparations for the budget for the financial year ended 30 June 2020, the accountant completed a breakeven analysis and concluded that the breakeven production and sales volume amounted to 20 500 units. The fact that the company only sold 20 000 units but is reporting a preliminary profit of R112 000 for the year ended 30 June 2020 has raised a concern about the integrity of the information generated by finance. You established the following:
1. The draft income statement for the year ended 30 June 2020 is as follows:
Amount (ZAR)
Sales
2 000 000
Cost of sales
1 628 000
o Inventory at beginning of year

o Raw materials
500 000
o Direct labour costs
600 000
o Manufacturing overheads
935 000
o Inventory at the end of the year: finished products
(407 000)
Gross profit
372 000
Fixed administration costs
80 000
Selling costs
180 000
Profit before tax
112 000
2. Sales and production had been budgeted for the 30 June 2020 financial year at 22 000 units. The budgeted selling price for the 30 June 2020 financial year was R100 per unit. Actual production volumes amounted to 25 000 units. The actual unit costs and selling prices as well as fixed costs were all equal to budgeted amounts. Labour is a variable cost. Total actual and budgeted fixed manufacturing overhead costs incurred amounted to R660 000.
3. The company has access to reliable supply of raw materials and therefore does not carry any raw materials inventory. There were no work-in-progress inventories at the beginning or end of the year. Manufacturing overheads and selling costs comprise both fixed and variable costs. Selling costs would have amounted to R190 000 at budgeted sales volumes of 22 000 units.

a)Reperform the calculation of the breakeven production and sales volumes based on the 30 June 2020 budget assumptions.

b)Discuss the reasons for the differences between the actual profit of R112 000 and the profit that would arise if the variable costing method was to be used.

c)Prepare the budgeted income statement (i.e. performance budget) of Slam (Pty) Ltd for the year ending 30 June 2021 presented using absorption costing method.

d)Advise the managing director on possible steps that could be taken to enable the company to achieve its target profit before tax of 10% of revenue.
Support your discussion with calculations.

e)Advise the managing director on possible steps that could be taken to enable the company to achieve its target profit before tax of 10% of revenue.
Support your discussion with calculations.
Answer #1

a) Calculation of Break even production and sales

BEP = Fixed cost / Contribution per unit

Contribution = Selling price - Variable cost

Fixed costs

Fixed Admin cost 80000
Selling cost 80000
Manufcaturing overheads fixed 660000
820000

Selling cost includes both variable and fixed elements as per the question. So we can use HI-LO method to find out the variable and fixed part from the selling costs

Hi 22000 units 190000
Lo 20000 units 180000
2000 units 10000

So Variable selling cost per unit = $10000/2000 = $5 per unit

Suppliment the value to Lo - (20000 units x $5)+ FC = $180000

FC = 180000-100000 = $80000

Variable cost

Rawmaterials 20
Labour cost 24
Manufacturing overhead 11
Selling cost 5
60

Total Manufacturing overheads incurred for producing 25000 units = 935000

Fixed production overhead (as budgeted) = 660000

Varaible overhead = 275000

Variable overhead per unit = 275000/25000 = $11 per unit

Contribution = SP - VC = 100 - 60 = $40 per unit

BEP = FC/Contribution per unit = 820000/40 = 20500 units.

Break even production and sales unit was amounted to 20500 units. There is no point in challenging the integrity of calculation of BEP.

b) The reasons for difference in profit

Even with a sale of 20000 units the firm made a profit despite the fact that the BEP was 20500 units.

It was only because they were able to get the advantage of large scale operation. Here they produced 25000 units which helps them to reduce the fixed cost per unit. This in turn changed to a profit.

c) Budgeted Income statement (Absorption Costing) (22000units)

$ $
Sales 2200000
Less cost
Rawmaterials 440000
Labour cost 528000
Manufacturing overhead 902000
1870000
1870000
Gross profit 330000
Less expenses
Selling cost 190000
Admin cost 80000 270000
Profit 60000

d) Target profit = 2000000 *10% = 200000

Number of units to be produced to get this tagret profit = (Fixed cost + target profit ) / Contribution per unit

= (820000+200000 )/ 40 = 25500 units.

They have to produce and sell 25500 units to get this target profit.